OpenStudy (carlyleukhardt):

What value of y makes the equation true? 3.4 + y = –6.85 A. –10.25 B. –3.25 C. 3.25 D. 10.25

OpenStudy (anonymous):

A

OpenStudy (carlyleukhardt):

Thank you <3

OpenStudy (anonymous):

Just plug in the values...

OpenStudy (mrnood):

@carlyleukhardt you cannot know if that is just a guess

OpenStudy (anonymous):

You're welcome

OpenStudy (carlyleukhardt):

DId she guess?

OpenStudy (mrnood):

@kenna_098 may just suck at maths! To solve this question: subtract 3.4 from each side of the equation and then you can simplify to get the answer

OpenStudy (carlyleukhardt):

Thats rude. But okay.

OpenStudy (anonymous):

I didnt guess... all you had to was subtract 3.4 from both sides It cancels out on the left side so on the right side of the equal sign you subtract -6.85-3.4 You end up getting -10.25 That means it is answer A That means that I am correct That also means that you are wrong because there is no simplifying involved!!! @MrNood

OpenStudy (carlyleukhardt):

Wow your smart

OpenStudy (carlyleukhardt):

But please guys don't fight

OpenStudy (mrnood):

The point I am making is that if you just give the answer then the poster has NO WAY of knowing whether you have any idea about how to do the question. MANY answers here are just plain wrong. you 'simplified when you completed the subtraction: 3.4+y= -6.85 3.4+y-3.4 = -6.85-3.4 SIMPLIFIES TO y= -10.25

OpenStudy (anonymous):

subtracting isn't simplification -10.25 is the answer you get from subtraction @MrNood

OpenStudy (mrnood):

giving the answer simply shows that YOU know how to do it - it doesn't help the poster to understand or learn anything

OpenStudy (anonymous):

I did show how I did it... In the first comment I showed my answer then in the next comment I explained how I got my answer @MrNood